OpenStudy (anonymous):

WRITE THE EXPRESSION AS A SINGLE LOGARITHM WITH THE COEFFICIENT IS 1. 10LOG(X+7)+3LOG(X-1)-7LOG(X-2)

OpenStudy (anonymous):

\[10*log{(x+1) + 3*log{(x-1)} - 7 *log{(x-2)}} \]

OpenStudy (anonymous):

we can review rules of logarithms, \[a*log{(b)} = log{(b^a)}\]

OpenStudy (anonymous):

so, let's apply this to each part of the equation: \[\log{((x+1)^{10})} +log{((x-1)^3}) - log{((x-2)^7)} \]

OpenStudy (anonymous):

and another rule: \[log(a) - log(b) = log(a/b)\]

OpenStudy (anonymous):

and \[log(a) + log(b) = log(a*b)\]

OpenStudy (anonymous):

so we can condense this: \[\log{((x+1)^{10})} +log{((x-1)^3}) - log{((x-2)^7)}\] using those rules

OpenStudy (anonymous):

would you like to give it a try?

OpenStudy (anonymous):

What did you get for the final answer?

OpenStudy (anonymous):

I haven't worked out that far

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

I GOT -4 THATS WRONG

OpenStudy (anonymous):

OK, just apply the rules of multiplying and dividing logarithms and you'll get the final answers, just 2 more steps are left

OpenStudy (anonymous):

I TRIED

OpenStudy (anonymous):

ok, let's look at the first two terms then

OpenStudy (anonymous):

because they are being added together, then we can use the multiplication rule

OpenStudy (anonymous):

of logarithms

OpenStudy (anonymous):

\[\log{((x+1)^{10})} +log{((x-1)^3})\] should now be of the form \[log(a*b)\]

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

\[log((x+1)^{10} * (x-1)^3)\]

OpenStudy (anonymous):

so you have now: \[log((x+1)^{10} * (x-1)^3) - log((x-2)^7)\]

OpenStudy (anonymous):

we apply the rule of division: \[log(a)-log(b) = log(a/b)\]

OpenStudy (anonymous):

\[log((x+1)^{10} * (x-1)^3) - log((x-2)^7)\] now is \[log(\frac{(x+1)^{10} * (x-1)^3}{(x-2)^7})\]

OpenStudy (anonymous):

there you have it

OpenStudy (anonymous):

so, you got -4?